### v1

# Gapfree graphs and powers of edge ideals with linear quotients

Abstract.
Let $G$ be a gapfree graph and let $I(G)$ be its edge ideal. An open conjecture of Nevo and Peeva states that $I(G)^{q}$ has a linear resolution for $q\gg 0$. We investigate a stronger conjecture that if $I(G)^{q}$ has linear quotients for some integer $q$, then $I(G)^{q+1}$ also has linear quotients. We give a partial solution to this conjecture. It is known that if $G$ does not contain a cricket, a diamond, or a $C_{4}$, then $I(G)^{q}$ has a linear resolution for $q\geq 2$. We construct a family of gapfree graphs $G$ containing cricket, diamond, $C_{4}$ and $C_{5}$ as induced subgraphs of $G$ for which $I(G)^{q}$ has linear quotients for $q\geq 2$.
Key words and phrases:
edge ideal, gapfree graph, linear quotients.
2020 Mathematics Subject Classification:
05E40, 13D02.
1. Introduction
Let $G$ be a finite simple graph, i.e., a graph that does not contain loops or multiple edges, with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. Let $S=K[x_{1},\ldots,x_{n}]$ denote the polynomial ring in $n$ variables over a field $K$. The edge ideal of $G$, denoted by $I(G)$, is the ideal of $S$ generated by squarefree quadratic monomials $x_{i}x_{j}$ for all $\{i,j\}\in E(G)$ (cf. [12]). The study of edge ideals has become a central topic in combinatorial commutative algebra, showcasing the deep interplay between combinatorial methods and algebraic structures. The present paper builds upon this rich foundation, advancing the understanding of powers of edge ideals and their algebraic properties.
A fundamental question in examining homogeneous ideals is when these ideals have linear resolution. In 2004, Herzog, Hibi and Zheng [9] showed that if $I(G)$ has linear resolution then all powers of $I(G)$ also have linear resolution. By a result of Fröberg [7], this condition is equivalent to the complementary graph of $G$ being chordal. A finite simple graph is called chordal if every cycle of length at least 4 has a chord, i.e., an edge connecting two non-consecutive vertices on the cycle. The proof of Herzog, Hibi and Zheng relies on Dirac’s theorem on chordal graphs, a cornerstone of classical graph theory, and leverages Gröbner bases techniques.
Furthermore, Francisco, Hà and Van Tuyl (see [11]) observed in unpublished work that if $I(G)^{q}$ has linear resolution for some $q\in{\mathbb{N}}$ then $G$ must be gapfree. A finite simple graph $G$ is gapfree if, for any two edges $e$ and $e^{\prime}$ with $e\cap e^{\prime}=\emptyset$, there is $f\in E(G)$ with $e\cap f\neq\emptyset$ and $e^{\prime}\cap f\neq\emptyset$. This led to the natural question:
“Is $G$ being gapfree both a necessary and sufficient condition for $I(G)^{q}$ to have linear resolution for all $q\geq 2$?”
In 2013, Nevo and Peeva [11] answered this question negatively by constructing a gapfree graph $G$ for which $I(G)^{2}$ does not have linear resolution. In light of this investigation, Nevo and Peeva [11] raised the following conjecture:
Conjecture 1.1.
If $G$ is gapfree, then $I(G)^{q}$ has linear resolution for all $q\gg 0$.
Conjecture 1.1 has garnered significant interest (cf. [2, 5, 6, 10] and references therein), but remains unresolved in general. While it has been verified for $q=2,3$ when $I(G)$ has small regularity (see [10]) and for random graphs with restricted edge probability (see [3]), progress has been limited. A promising approach to this challenging conjecture is to explore the stronger property of having linear quotients. Specifically, we make the following conjecture:
Conjecture 1.2.
If $G$ is gapfree and $I(G)^{q}$ has linear quotients for some $q\in{\mathbb{N}}$, then $I(G)^{N}$ has linear quotients for all $N\geq q$.
For a subset $W\subset V(G)$, the induced subgraph $G_{W}$ is the graph with vertex set $W$ consisting of all edges $\{i,j\}\in E(G)$ where $i,j\in W$. It was shown in [1, 5, 6] that $I(G)^{q}$ have linear resolution for all $q\geq 2$ if $G$ is gapfree and contains no induced subgraph isomorphic to: (i) cricket; (ii) diamond; (iii) $C_{4}$ (see Figure 1).
Figure 1. Cricket; Diamond; $C_{4}$
For instance, the pentagon $C_{5}$ is gapfree, but its complementary graph is not chordal. Consequently, $I(C_{5})$ does not have a linear resolution. However, $C_{5}$ does not contain any induced subgraph that is isomorphic to a cricket (or a diamond, or $C_{4}$), so $I(C_{5})^{q}$ have linear resolution for all $q\geq 2$. On the other hand, it was shown (cf. [8, Corollary 10.1.7 and Theorems 10.1.9 and 10.2.5]) that if the complementary graph of $G$ is chordal then all powers of $I(G)$ have linear quotients. Beside the anticycles, as exhibited in a recent preprint [4] which appears after the first draft of this paper was written, there are currently no other known examples of a gapfree graph $G$ whose complementary graph is not chordal, yet for which all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. This motivates the search for natural classes of such graphs.
In this paper, we establish a framework for addressing Conjecture 1.2. In Section 2, we examine the duplication technique which uses duplication of vertices in $G$ to produce new graphs $G^{*}$. We show that if a power $I(G)^{q}$ has linear quotients, then so does $I(G^{*})^{q}$ for certain derived graphs $G^{*}$ (Proposition 2.2). In Section 5, a similar result is shown when the vertex duplication is replaced by a clique, as long as the resulting graph is still gapfree (Theorem 5.3).
Section 3 studies Conjecture 3.1, which makes a strong case for Conjecture 1.2. Particularly, Proposition 3.2 shows that Conjecture 3.1 holds for the pentagon $C_{5}$. In other words, all powers $I(C_{5})^{q}$ with $q\geq 2$ have linear quotients. Section 4 extends the study in Section 3 to a special class of graphs which are gapfree but may contain critical induced subgraphs such as crickets, diamonds and $C_{4}$. In Propositions 4.2 and 4.3, it is shown that, for these graphs, all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. The main result of the section (Theorem 4.6) extends the aforementioned result to an infinite family of graphs.
Finally, Section 6 presents a refined analysis of modified version of Conjectures 3.1 and 1.2, culminating in Theorem 6.4, which exhibits that if $I(G)^{2},\ldots,I(G)^{7}$ have linear quotients with a special ordering of their generators, then $I(G)^{q}$, for all $q\geq 2$, have linear quotients. This result reduces the verification of Conjectures 3.1 and 1.2 to finitely many powers and, therefore, suggests that resolving these conjectures may be within reach.
2. Duplication of finite graphs
Let $G$ be a finite graph on $V(G)=\{1,\ldots,n\}$ with $E(G)=\{e_{1},\ldots,e_{s}\}$ and $S=K[x_{1},\ldots,x_{n}]$ the polynomial ring in $n$ variables over a field $K$. For an edge $e=\{i,j\}\in E(G)$, we define $u_{e}=x_{i}x_{j}\in S$. Recall that the edge ideal of $G$ is the ideal $I(G)$ of $S$ which is generated by those monomials $u_{e}$ with $e\in E(G)$. To simplify the notation, unless misleading, we may write $e$ instead of $u_{e}$. For example, $e_{1}^{3},e_{1}e_{2}^{2}\in I(G)^{3}$.
Let, in general, $I$ be an ideal of $S$ which is generated by monomials of the same degree and $\{u_{1},\ldots,u_{r}\}$ the unique minimal system of monomial generators of $I$. We then call $u_{1},\ldots,u_{r}$ the generators of $I$. We say that $I$ has linear quotients if there is an ordering $u_{1}>\cdots>u_{r}$ of the generators of $I$ for which the colon ideal $(u_{1},\ldots,u_{i}):u_{i+1}$ is generated by a subset of $\{x_{1},\ldots,x_{n}\}$ for each $1\leq i<r$. If $I$ has linear quotients, then $I$ has linear resolution ([8, Proposition 8.2.1]).
Definition 2.1.
Let $G$ be a finite graph and fix a vertex $a$ of $G$. Let $G^{*}$ be a finite graph obtained by adding a new vertex $a^{*}$ to $G$ with
$E(G^{*})=E(G)\cup\{a^{*}b:ab\in E(G)\}.$
We say that $G^{*}$ is obtained from $G$ by duplicating a vertex $a$ of $G$. Note that $aa^{*}\not\in E(G^{*})$.
Proposition 2.2.
If $I(G)^{s}$ has linear quotients, then $I(G^{*})^{s}$ has linear quotients.
Proof.
Let $H=G^{*}$ be the graph obtained from $G$ by duplicating the vertex $x\in V(G)$. Let $y\in V(H)$ be the duplicate of $x$ in $H$. Let $\mathcal{G}(I)$ denote the minimal generating set of $I$. Let $u\in\mathcal{G}(I(G)^{s})$ with $x$-degree $r>0$. Then $u$ will produce $r$ new generators in $\mathcal{G}(I(H)^{s})$, namely
(1) $\frac{\,u\,}{x}y,\frac{\,u\,}{x^{2}}y^{2},\dots,\frac{\,u\,}{x^{r}}y^{r}$
In fact, all of the new generators will be of this form since $x$ and $y$ are non-adjacent in $H$. Extend the linear quotients order on $\mathcal{G}(I(G)^{s})$ to a linear quotients order on $\mathcal{G}(I(H)^{s})$ as follows: If $u\in\mathcal{G}(I(G)^{s})$ has positive $x$-degree $r$, then insert the new generators as in (1) right after $u$.
Let $w_{1},\dots,w_{q}$ be the order on the minimal generators of $I(H)^{s}$. Let $t\geq 2$. We will show that $I=(w_{1},\dots,w_{t-1}):w_{t}$ is generated by variables. Let $\ell<t$ be fixed. We will show that there is a variable generator of $I$ that divides $w_{\ell}:w_{t}$.
(CASE 1) Suppose that $y$ does not divide $w_{t}$. Let $J$ be the ideal generated by $\{w_{i}:w_{t}\mid 1\leq i<t,w_{i}\in\mathcal{G}(I(G)^{s})\}$. Then $J$ is generated by variables and $\mathcal{G}(J)\subseteq\mathcal{G}(I)$. Therefore, we may assume that $y$ divides $w_{\ell}$.
We now claim that $x\in\mathcal{G}(J)$ implies $y\in\mathcal{G}(I)$. Indeed, if $x\in\mathcal{G}(J)$, then $w_{t^{\prime}}:w_{t}=x$ for some $w_{t^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $t^{\prime}<t$. Since $w_{t^{\prime}}$ has positive $x$-degree, by definition of the order, $w_{t^{\prime\prime}}=w_{t^{\prime}}y/x$ for some $t^{\prime}<t^{\prime\prime}<t$. Then $w_{t^{\prime\prime}}:w_{t}=y$ and $y\in\mathcal{G}(I)$ as wanted.
Case 1.1: If $x\in\mathcal{G}(J)$, then by the above claim $y\in\mathcal{G}(I)$, and thus $y$ divides $w_{\ell}:w_{t}$.
Case 1.2: Suppose that $x\notin\mathcal{G}(J)$. Let $w_{\ell}=x^{q_{1}}y^{q_{2}}v$ where $q_{1}\geq 0$ is the $x$-degree of $w_{\ell}$ and $q_{2}\geq 1$ is the $y$-degree of $w_{\ell}$. By definition of the order, $w_{\ell^{\prime}}=x^{q_{1}+q_{2}}v$ for some $\ell^{\prime}<\ell$. Then, there is a variable generator $z\neq x$ of $J$ such that $z$ divides $w_{\ell^{\prime}}:w_{t}$. This implies that $z$ divides $v:w_{t}$. Thus, $z$ divides $w_{\ell}:w_{t}$ as desired.
(CASE 2) Suppose that $y$ divides $w_{t}$. First, we claim that $x\in\mathcal{G}(I)$. Indeed, by definition of the order, $w_{\eta}=xw_{t}/y$ for some $\eta<t$ and thus $w_{\eta}:w_{t}=x\in\mathcal{G}(I)$.
Let $w_{\ell}=x^{\gamma}y^{\sigma}N$ where $\gamma$ is the $x$-degree, and $\sigma$ is the $y$-degree of $w_{\ell}$. Let $w_{t}=x^{\alpha}y^{\beta}M$ where $\alpha$ is the $x$-degree, and $\beta\geq 1$ is the $y$-degree of $w_{t}$. Let $w_{L}=x^{\gamma+\sigma}N$ and $w_{T}=x^{\alpha+\beta}M$. Then, by definition of the order, $L<\ell<T<t$.
If $\gamma>\alpha$, then nothing is left to show as $x\in\mathcal{G}(I)$. Therefore, let us assume that $\gamma\leq\alpha$.
Case 2.1: Suppose that $\gamma+\sigma\leq\alpha+\beta$. Since $I(G)^{s}$ has linear quotients, there exists $w_{T^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $T^{\prime}<T$ such that $w_{T^{\prime}}:w_{T}=z$ for some variable $z$ that divides $w_{L}:w_{T}=N:M$. Then $z$ divides $w_{\ell}:w_{t}$. Note that $z\neq x,y$. It remains to show that $z$ is a generator of $I$. Since $w_{T}$ and $w_{T^{\prime}}$ are of the same degree, and $w_{T^{\prime}}:w_{T}=z$, either
  • (a)
    $w_{T^{\prime}}=x^{\alpha+\beta-1}Mz$ or
  • (b)
    $w_{T^{\prime}}=x^{\alpha+\beta}Mz/z^{\prime}$ for some variable $z^{\prime}(\neq z)$ dividing $M$.
If (a) holds, then $w_{t^{\prime}}=x^{\alpha}y^{\beta-1}Mz$ for some $t^{\prime}$ that satisfies $T^{\prime}\leq t^{\prime}<T<t$. Since $w_{t^{\prime}}:w_{t}=z$, it follows that $z\in\mathcal{G}(I)$ as wanted.
If (b) holds, then $w_{t^{\prime}}=x^{\alpha}y^{\beta}Mz/z^{\prime}$ for some $t^{\prime}$ that satisfies $T^{\prime}<t^{\prime}<T<t$. Since $w_{t^{\prime}}:w_{t}=z$, it follows that $z\in\mathcal{G}(I)$ as wanted.
Case 2.2: Suppose that $\gamma+\sigma>\alpha+\beta$. Since we are also assuming that $\gamma\leq\alpha$, it follows that $\sigma>\beta$. Then $w_{\ell}:w_{t}=y^{\sigma-\beta}(N:M)$. Then, it remains to show that either $y\in\mathcal{G}(I)$, or there is a variable in $\mathcal{G}(I)$ that divides $N:M$.
Since $I(G)^{s}$ has linear quotients, there exists $w_{L^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $L^{\prime}<T$ such that $w_{L^{\prime}}:w_{T}=z$ for some variable $z$ that divides $w_{L}:w_{T}=x^{\gamma+\sigma-(\alpha+\beta)}(N:M)$.
Case 2.2.1: Suppose that $z=x$. Then, $w_{L^{\prime}}=x^{\alpha+\beta+1}M/z^{\prime}$ for some variable $z^{\prime}$ dividing $M$. Then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta+1}M/z^{\prime}$ for some $\ell^{\prime}$ that satisfies $L^{\prime}<\ell^{\prime}<T<t$. It follows that $w_{\ell^{\prime}}:w_{t}=y$ and $y\in\mathcal{G}(I)$ as desired.
Case 2.2.2: Suppose that $z\neq x$, and $z$ divides $N:M$. It suffices to show that $z\in\mathcal{G}(I)$. Since $w_{L^{\prime}}$ and $w_{T}$ have the same degree, either
  • (a)
    $w_{L^{\prime}}=x^{\alpha+\beta-1}Mz$ or
  • (b)
    $w_{L^{\prime}}=x^{\alpha+\beta}Mz/z^{\prime}$ for some variable $z^{\prime}(\neq z)$ dividing $M$.
If $(a)$ holds, then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta-1}Mz$ for some $\ell^{\prime}$ with $L^{\prime}\leq\ell^{\prime}<T<t$. Then, $w_{\ell^{\prime}}:w_{t}=z$ and $z\in\mathcal{G}(I)$ as desired.
If $(b)$ holds, then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta}Mz/z^{\prime}$ for some $\ell^{\prime}$ with $L^{\prime}<\ell^{\prime}<T<t$. Then, $w_{\ell^{\prime}}:w_{t}=z$ and $z\in\mathcal{G}(I)$ as desired.
3. Ordering of pure powers
Let $G$ be a gapfree graph with $E(G)=\{e_{1},\ldots,e_{s}\}$ and suppose that a power $I(G)^{q}$ of the edge ideal $I(G)$ has linear quotients with an ordering
${\mathcal{N}}:u_{1}>\cdots>u_{r}$
of the generators of $I(G)^{q}$, where pure powers appear to be
$e_{1}^{q}>\cdots>e_{s}^{q}.$
Then the ordering
${\mathcal{N}}^{(1)}:u_{1}e_{1}>\cdots>u_{r}e_{1}>u_{1}e_{2}>\cdots>u_{r}e_{2}>\cdots>u_{r}e_{s}$
of the generators of $I(G)^{q+1}$, where $u_{i}e_{j}$ is omitted if $u_{i}e_{j}=u_{k}e_{\ell}$ with $\ell<j$, is called the efficient ordering coming from the ordering ${\mathcal{N}}$. Recurrently we define the ordering ${\mathcal{N}}^{(j)}$ of the generators of $I(G)^{q+j}$ for $j\geq 2$ by setting
${\mathcal{N}}^{(j)}=({\mathcal{N}}^{(j-1)})^{(1)}.$
Strong evidence for Conjecture 1.2 is the following:
Conjecture 3.1.
Let $G$ be a gapfree graph with $E(G)=\{e_{1},\ldots,e_{s}\}$ for which a power of $I(G)$ has linear quotients. Let $q_{0}$ denote the smallest integer for which $I(G)^{q_{0}}$ has linear quotients. Suppose that $I(G)^{q_{0}}$ has linear quotients with an ordering ${\mathcal{N}}:u_{1}>\cdots>u_{r}$ of the generators of $I(G)^{q_{0}}$. Then each power $I(G)^{q}$ for $q>q_{0}$ has linear quotients with the ordering ${\mathcal{N}}^{(q-q_{0})}$ of the generators of $I(G)^{q}$.
Let $C_{5}$ be the pentagon on the vertex set $\{a,b,c,d,e\}$ with the edges
$e_{1}=ab,\,\,e_{2}=bc,\,\,e_{3}=cd,\,\,e_{4}=de,\,\,e_{5}=ea.$
One can easily prove directly that $I(C_{5})^{2}$ has linear quotient with the following ordering, with omitting $>$, of its $15$ generators:
$a^{2}b^{2},ab^{2}c,b^{2}c^{2},abcd,bc^{2}d,c^{2}d^{2},abce,a^{2}be,abde,bcde,cd^{2}e,acde,ade^{2},a^{2}e^{2},d^{2}e^{2}.$
By using $e_{1},e_{2},e_{3},e_{4},e_{5}$, the above ordering is
(2) $e_{1}^{2},e_{1}e_{2},e_{2}^{2},e_{1}e_{3},e_{2}e_{3},e_{3}^{2},e_{2}e_{5},e_{1}e_{5},e_{1}e_{4},e_{2}e_{4},e_{3}e_{4},e_{3}e_{5},e_{4}e_{5},e_{5}^{2},e_{4}^{2}.$
The ordering of pure powers is
$e_{1}^{2}>e_{2}^{2}>e_{3}^{2}>e_{5}^{2}>e_{4}^{2}.$
We show that Conjecture 3.1 is true for the pentagon with $q_{0}=2$.
Proposition 3.2.
All powers $I(C_{5})^{q}$ with $q\geq 2$ have linear quotients.
Proof.
Starting from the linear quotient ordering (2) of $I(C_{5})^{2}$, we prove that Conjecture 3.1 is true for $q\geq 2$ by using induction on $q$. Note that $e_{1}=ab,e_{2}=bc,e_{3}=cd,e_{4}=de$ and $e_{5}=ea$ are algebraically independent in $S=K[a,b,c,d,e]$.
Let $q>3$ and suppose that $I(C_{5})^{q-1}$ has linear quotients with the ordering $u_{1}>\cdots>u_{\binom{5+(q-1)-1}{q-1}}$ coming from Conjecture 3.1. In particular $u_{1}=e_{1}^{q-1}$ and $u_{\binom{5+(q-1)-1}{q-1}}=e_{4}^{q-1}$. Let $w_{1}>\cdots>w_{\binom{5+q-1}{q}}$ denote the ordering of $I(C_{5})^{q}$ coming from Conjecture 3.1.
Let ${\mathcal{A}}(e_{1})$ denote the set of those $w_{i}$ which is divided by $e_{1}$, ${\mathcal{A}}(e_{2})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})$ which is divided by $e_{2}$, ${\mathcal{A}}(e_{3})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$ which is divided by $e_{3}$ and ${\mathcal{A}}(e_{5})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$ which is divided by $e_{5}$. Finally, ${\mathcal{A}}(e_{4})=\{e_{4}^{q}\}$.
We claim that the colon ideal $(w_{1},\ldots,w_{p-1}):w_{p}$ is generated by a subset of $\{a,b,c,d,e\}$. If $w_{p}\in{\mathcal{A}}(e_{1})$, then the claim follows from the induction hypothesis.
Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ with $i<p$. Let $w_{i}=u_{s}e_{2}$ and $w_{p}=u_{t}e_{2}$ with $s<t$. Then there is $u_{\ell}$ with $\ell<t$ for which $u_{\ell}:u_{t}$ is a variable and $(u_{\ell}:u_{t})|(u_{s}:u_{t})$. One has $j<p$ with $w_{j}=u_{\ell}e_{2}$. Note that $w_{j}$ may belong to ${\mathcal{A}}(e_{1})$. Then $w_{j}:w_{p}$ is a variable and $(w_{j}:w_{p})|(w_{i}:w_{p})$, as desired.
Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{1})$. If $a$ does not appear in $w_{p}$, then $w_{i}:w_{p}$ is divided by $a=(e_{1}w_{p}/e_{2}):w_{p}$. If $e_{5}=ae$ divides $w_{p}$, then both $(e_{1}w_{p}/e_{5}):w_{p}=b$ and $(e_{1}e_{3}w_{p}/e_{2}e_{5}):w_{p}=d$ belong to $(w_{1},\ldots,w_{p-1}):w_{p}$. Let $w_{i}=e_{1}^{f_{1}}e_{2}^{f_{2}}e_{3}^{f_{3}}e_{4}^{f_{4}}e_{5}^{f_{5}}$ and $w_{p}=e_{2}^{f^{\prime}_{2}}e_{3}^{f^{\prime}_{3}}e_{4}^{f^{\prime}_{4}}e_{5}^{f^{\prime}_{5}}$. If one of the inequalities
$f_{1}+f_{5}>f^{\prime}_{5},\,\,\,f_{1}+f_{2}>f_{2}^{\prime},\,\,\,f_{3}+f_{4}>f^{\prime}_{3}+f^{\prime}_{4}$
holds, then $w_{i}:w_{p}\in(a,b,d)$. If none of the above inequalities is satisfied, then
$2f_{1}+f_{2}+f_{3}+f_{4}+f_{5}\leq f_{2}^{\prime}+f_{3}^{\prime}+f^{\prime}_{4}+f^{\prime}_{5},$
which is impossible, since $f_{1}>0$.
Let $w_{p}\in{\mathcal{A}}(e_{3})$. Since $e_{2}(w_{p}/e_{3})\in{\mathcal{A}}(e_{2})$, it follows that $(e_{2}w_{p}/e_{3}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $e_{i}\in{\mathcal{A}}(e_{3})$, then the similar argument as in the case of $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ is valid without modification.
Let $w_{p}\in{\mathcal{A}}(e_{5})$. Since $e_{1}(w_{p}/e_{5})\in{\mathcal{A}}(e_{1})$, it follows that $(e_{1}w_{p}/e_{5}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $e_{4}$ divides $w_{p}$, then $e_{3}(w_{p}/e_{4})\in{\mathcal{A}}(e_{3})$ and $(e_{3}w_{p}/e_{4}):w_{p}=c$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $c$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $c$ divides $w_{i}:w_{p}$. Let $w_{p}=e_{5}^{q}$. Since $e_{4}e_{5}^{q-1}>e_{5}^{q}$ and $e_{4}e_{5}^{q-1}:e_{5}^{q}=d$ belongs $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $d$ does not appear in $w_{p}=e_{5}^{q}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $d$ divides $w_{i}:w_{p}$. If $e_{i}\in{\mathcal{A}}(e_{5})$, then the similar argument as in the case of $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ is valid without modification.
Finally, if $w_{p}\in{\mathcal{A}}(e_{4})$, in other words, $e_{p}=e_{4}^{q}$, then a routine computation done as above yields $(w_{1},\ldots,w_{p-1}):w_{p}=(a,c)$. This completes the proof.
Remark 3.3.
The following ordering
$a^{2}b^{2},a^{2}be,ab^{2}c,abce,a^{2}e^{2},b^{2}c^{2},abcd,abde,bc^{2}d,ade^{2},bcde,acde,cd^{2}e,c^{2}d^{2},d^{2}e^{2}.$
can also be an ordering for $I(C_{5})^{2}$ to have linear quotients, where the ordering of pure powers is $e_{1}^{2}>e_{5}^{2}>e_{2}^{2}>e_{3}^{2}>e_{4}^{2}$.
4. Pentagon together with vertices
Developing the argument done in the proof of Proposition 3.2, we now study finite graphs formed by carefully adding vertices to the pentagon. We first study the case where a single vertex is added with specific connecting edges.
Let $G$ be the finite graph with the vertices $\{a,b,p,q,x,z\}$ and with edges
$e_{1}=\{a,b\},e_{2}=\{a,x\},e_{3}=\{b,x\},e_{4}=\{a,p\},e_{5}=\{b,q\},$
$e_{6}=\{p,z\},e_{7}=\{x,z\},e_{8}=\{q,z\}$
depicted below (Figure $2$).
$z$$p$$q$$x$$a$$b$
Figure 2. Pentagon together with one vertex
The edge ideal $I(G)$ has $8$ generators and, since $e_{2}e_{6}=e_{4}e_{7}$ and $e_{3}e_{8}=e_{5}e_{7}$, it follows that $I(G)^{2}$ has $\binom{8+1}{2}-2=36-2=34$ generators. Let ${\mathcal{N}}$ denote the following order these generators:
$\begin{array}[]{cccc}e_{1}^{2}=(ab)^{2},&e_{1}e_{2}=(ab)(ax),&e_{1}e_{3}=(ab)(bx),&e_{1}e_{4}=(ab)(ap),\\ e_{1}e_{5}=(ab)(bq),&e_{2}e_{4}=(ax)(ap),&e_{2}e_{3}=(ax)(bx),&e_{3}e_{5}=(bx)(bq),\\ e_{3}e_{4}=(bx)(ap)&e_{2}e_{5}=(ax)(bq),&e_{4}e_{5}=(ap)(bq),&e_{1}e_{6}=(ab)(pz),\\ e_{1}e_{8}=(ab)(qz),&e_{1}e_{7}=(ab)(xz),&e_{4}e_{7}=(ap)(xz),&e_{5}e_{7}=(bq)(xz),\\ e_{3}e_{6}=(bx)(pz),&e_{3}e_{8}=(ax)(qz),&e_{2}e_{7}=(ax)(xz)&e_{3}e_{7}=(bx)(xz),\\ e_{4}e_{6}=(ap)(pz),&e_{5}e_{8}=(bq)(qz),&e_{4}e_{8}=(ap)(qz),&e_{5}e_{6}=(bq)(pz),\\ e_{6}e_{7}=(pz)(xz),&e_{6}e_{8}=(pz)(qz),&e_{7}e_{8}=(xz)(qz),&e_{4}^{2}=(ap)^{2},\\ e_{2}^{2}=(ax)^{2},&e_{3}^{2}=(bx)^{2},&e_{5}^{2}=(bq)^{2},&e_{6}^{2}=(pz)^{2},\\ e_{7}^{2}=(xz)^{2},&e_{8}^{2}=(qz)^{2}.&&\\ \end{array}$
It is now a direct computation to verify that with the order ${\mathcal{N}}$, $I(G)^{2}$ has linear quotients.
Using the ordering ${\mathcal{N}}$ of the generators of $I(G)^{2}$ as above, one can inductively show that all powers of $I(G)$ have linear quotients. The following notation will be useful in the proof of Proposition 4.2.
Definition 4.1.
Let $G$ be a finite graph and $w$ one of the generators of $I(G)^{t}$. The notation $e_{j}\shortmid w$ means that there exists a factorization $w=e_{i_{1}}\cdots e_{i_{t}}$ with $e_{j}=e_{i_{k}}$ for some $1\leq k\leq t$.
Note that this notation is intended to convey a notion of division in terms of a factorization into a product of generators, rather than literal division. For instance, if $G$ is as in Figure 2, then $e_{1}=ab$ divides $w=e_{4}e_{5}=abpq$, but $e_{1}\nshortmid w$ since $pq$ is not a generator of $I(G)$.
Proposition 4.2.
Let $G$ be the finite graph of Figure $2$. Then all powers $I(G)^{N}$ with $N\geq 2$ have linear quotients.
Proof.
We prove that Conjecture 3.1 is true for $G$ with $q_{0}=2$. Assume $N\geq 3$ and suppose that $I(G)^{N-1}$ has linear quotients with the ordering ${\mathcal{N}}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{N-1}$ and $u_{r}=e_{8}^{N-1}$. Let ${\mathcal{N}}^{(1)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{N}$ coming from ${\mathcal{N}}$ as in Conjecture 3.1. For ease of notation, use the following notation to refer to elements in ${\mathcal{N}}^{(1)}$
${\mathcal{U}}e_{1}={\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8})$
where $w_{k}\in{\mathcal{A}}(e_{i})$ means $e_{i}\shortmid w_{k}$ and $e_{j}\nshortmid w_{k}$ for each $j<i$.
We claim that the colon ideal $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear for each $w_{n}$. If $w_{n}\in{\mathcal{A}}(e_{1})$, then the claim follows from the induction hypothesis.
Now, let $w_{n}=u_{s}e_{4}\in{\mathcal{A}}(e_{4})$ for $s\leq r$. First, observe that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{4})=(b)$ for $u_{s}e_{1}\in{\mathcal{A}}(e_{1})$. Let $w_{i}=u_{t}e_{4}\in{\mathcal{A}}(e_{4})$ for $t<s$. By the induction hypothesis, there exists $u_{l}$ with $l<s$ such that $u_{l}:u_{s}$ is a variable dividing $u_{t}:u_{s}$. Let $w_{j}=u_{l}e_{4}$ for $j<n$. Then, $w_{j}:w_{n}$ is a variable and it divides $w_{i}:w_{n}$, as desired. Note that $w_{j}\in{\mathcal{A}}(e_{1})$ is possible.
Next, let $w_{i}\in{\mathcal{A}}(e_{1})$ and consider whether $b$ divides $w_{n}=u_{s}e_{4}$. Recall that $b\nmid e_{4}$. If $b\nmid u_{s}$, then $w_{i}:w_{n}\in(b)$ since every element of ${\mathcal{A}}(e_{1})$ is divisible by $b$. Then, $(u_{1}e_{1},\ldots,u_{r}e_{1}):w_{n}=(b)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case. Suppose $b\mid u_{s}$. Then either $e_{3}\shortmid u_{s}$ or $e_{5}\shortmid u_{s}$. If $e_{3}\shortmid u_{s}$, then $u_{s}=w^{\prime}e_{3}$ for a generator $w^{\prime}\in I(G)^{N-2}$. Consider the following elements $w_{j}=w^{\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. If $e_{5}\shortmid u_{s}$, then $u_{s}=w^{\prime\prime}e_{5}$ for a generators $w^{\prime\prime}\in I(G)^{N-2}$. Similarly, consider $w_{j}=w^{\prime\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, in either of these scenarios, we have $(a,b,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$ Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}\cdots e_{8}^{t_{8}^{\prime}}$ since $e_{1}\nshortmid w_{n}$. If $w_{i}:w_{n}\notin(a,b,z)$, then examining the degrees of $a,b,z$ in $w_{i}$ and $w_{n}$, respectively, results with
$t_{1}+t_{2}+t_{4}$$\leq$ $t_{2}^{\prime}+t_{4}^{\prime}$
$t_{1}+t_{3}+t_{5}$$\leq$ $t_{3}^{\prime}+t_{5}^{\prime}$
$t_{6}+t_{7}+t_{8}$$\leq$ $t_{6}^{\prime}+t_{7}^{\prime}+t_{8}^{\prime}$
which implies that $t_{1}\leq 0$, a contradiction. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear for each $w_{n}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})$.
Let $w_{n}=u_{s}e_{2}\in{\mathcal{A}}(e_{2})$ for some $s\leq r$. Note that $e_{6}\nshortmid w_{n}$ because $e_{6}e_{2}=e_{7}e_{4}$ and $e_{4}\nshortmid e_{n}$. Next, observe that $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{2})=(b)$ and $(u_{s}e_{4}:u_{s}e_{2})=(p)$. Let $w_{i}\in{\mathcal{A}}(e_{2})$. As in the previous case, by the induction hypothesis, there exists $w_{j}$ for $j<n$ such that $w_{j}:w_{n}$ is a variable and it divides $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})$. Observe that $p\nmid u_{s}$; otherwise, $e_{4}\shortmid u_{s}$ or $e_{6}\shortmid u_{s}$ but neither of these cases are possible since $w_{n}\in{\mathcal{A}}(e_{2})$. If $b\nmid u_{s}$, then $(w_{i}:w_{n})\subseteq(b,p)$ which implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is generated by variables. So, claim holds for this case. Suppose $b\mid u_{s}.$ Then either $e_{3}\shortmid u_{s}$ or $e_{5}\shortmid u_{s}$ since $e_{1}\nshortmid u_{s}.$ If $e_{3}\shortmid u_{s}$, then $u_{s}=w^{\prime}e_{3}=w^{\prime}bx$ for some $w^{\prime}\in I(G)^{N-2}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, $(a,b,p,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
Assuming $e_{5}\shortmid u_{s}$ also yields to the above containment. Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ or $t_{4}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}e_{3}^{t_{3}^{\prime}}e_{5}^{t_{5}^{\prime}}e_{t}^{t_{7}^{\prime}}e_{8}^{t_{8}^{\prime}}$ since $e_{1},e_{4},e_{6}$ do not divide $w_{n}=u_{s}e_{2}.$ If $w_{i}:w_{n}\notin(a,b,p,z)$, then examining degrees of $a,b,p,z$ in $w_{i}$ and $w_{n}$ results with $t_{1}+t_{4}+t_{6}\leq 0$, a contradiction. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
Now, let $w_{n}=u_{s}e_{3}\in{\mathcal{A}}(e_{3})$ for some $u_{s}\in I(G)^{N-1}$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{3})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})$. Note that this collection contains all generators of $I(G)^{N}$ that are divisible by $a$ since $e_{1},e_{2}$ and $e_{4}$ are the only edges incident to $a$ and $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
Next, let $w_{n}=u_{s}e_{5}\in{\mathcal{A}}(e_{5})$ for some $u_{s}\in I(G)^{N-1}$. First observe that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{5})=(a)$ and $(u_{s}e_{3}:u_{s}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that $w_{i}$ is divisible by $a$ or $x$ and $w_{n}$ is not divisible by $a$. Furthermore, $w_{n}$ is not divisible by $x$. Otherwise, if $x\mid u_{s}$, then $u_{s}=w^{\prime}e_{7}$ for a generator $w^{\prime}$ of $I(G)^{N-2}$ and $w_{n}=w^{\prime}e_{7}e_{5}=w^{\prime}e_{6}e_{3}$ which implies that $e_{3}\shortmid w_{n}$, a contradiction. So, $w_{i}:w_{n}\in(a,x)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
Let $w_{n}=u_{s}e_{6}\in{\mathcal{A}}(e_{6})$ for some $u_{s}\in I(G)^{N-1}$. Observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{4}:u_{s}e_{6})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{6})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})$. Note that $w_{i}$ is divisible by $a$ or $b$ while $a\nshortmid w_{n}$ and $b\nshortmid w_{n}$. This implies that $w_{i}:w_{n}\in(a,b)$.
Observe that $u_{s}=e_{6}^{t_{6}}e_{7}^{t_{7}}e_{8}^{t_{8}}$ since $u_{s}$ must be divisible by $z$. If $t_{8}\geq 1$, then $u_{s}=w^{\prime}e_{8}$ for a generator $w^{\prime}\in I(G)^{N-2}$ and $(w^{\prime}e_{6}e_{5}:w^{\prime}e_{6}e_{8})=(b)$. Similarly, if $t_{7}\geq 1$, then $u_{s}=w^{\prime}e_{7}$ for a generator $w^{\prime}\in I(G)^{q-2}$ and $(w^{\prime}e_{6}e_{3}:w^{\prime}e_{6}e_{7})=(b)$. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in these cases. For the remaining case, we have $w_{n}=e_{6}^{q}$. Notice that, for each $w_{i}$, we have $w_{i}:w_{n}\in(a,x,q)$ since $\{a,x,q\}$ is a vertex cover of $G\setminus\{e_{6}\}$. Furthermore, note that the generators $w_{j}=e_{6}^{N-1}e_{7}$ and $w_{k}=e_{6}^{N-1}e_{8}$ come before $w_{n}$ in ${\mathcal{A}}(e_{6})$ from the construction of ${\mathcal{N}}$. In addition, $(w_{j}:w_{n})=(x)$ and $(w_{j}:w_{n})=(q)$. Thus, the claim holds.
For the final two cases, if $w_{n}=u_{s}e_{7}\in{\mathcal{A}}(e_{7})$ then none of $a,b,p$ divide $u_{s},$ but every element of ${\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})\cup{\mathcal{A}}(e_{6})$ is divisible by one of these three variables. Since $(u_{s}e_{2}:u_{s}e_{7})=(a)$ and $(u_{s}e_{3}:u_{s}e_{7})=(b)$ and $(u_{s}e_{6}:u_{s}e_{7})=(p)$, we are done. For the last case, we have $w_{n}=e_{8}^{q}$. Note that $(e_{8}^{q-1}e_{5}:e_{8}^{q})=(b),~{}~{}(e_{8}^{N-1}e_{6}:e_{8}^{N})=(p)$ and $(e_{8}^{N-1}e_{7}:e_{8}^{N})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{8}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$. Therefore, the claim holds.
We now consider a different class of gapfree graphs built from adding a vertex to a pentagon.
$z$$p$$q$$x$$a$$b$
Figure 3. Pentagon together with one vertex and multiple inner edges
Proposition 4.3.
Let $G$ be the finite gapfree graph given in Figure 3. Then $I(G)^{N}$ has linear quotients for all $N\geq 2$.
Proof.
Label the edges of $G$ as follows:
$e_{1}=\{a,b\},~{}~{}e_{2}=\{a,p\},~{}~{}e_{3}=\{a,x\},~{}~{}e_{4}=\{b,x\},~{}~{}e_{5}=\{b,q\},$
$e_{6}=\{x,p\},~{}~{}e_{7}=\{x,q\},~{}~{}e_{8}=\{p,z\},~{}~{}e_{9}=\{q,z\}$
One can verify that $I(G)^{2}$ has linear quotients with respect to the following order:
$\begin{array}[]{cccc}e_{1}^{2}=(ab)^{2},&e_{1}e_{3}=(ab)(ax),&e_{1}e_{4}=(ab)(bx),&e_{1}e_{2}=(ab)(ap),\\ e_{1}e_{6}=(ab)(xp),&e_{1}e_{5}=(ab)(bq),&e_{1}e_{7}=(ab)(xq),&e_{1}e_{8}=(ab)(pz),\\ e_{1}e_{9}=(ab)(qz),&e_{2}e_{8}=(ap)(pz),&e_{2}e_{3}=(ap)(ax),&e_{2}e_{7}=(ap)(xq),\\ e_{2}e_{9}=(ap)(qz),&e_{2}e_{5}=(ap)(bq),&e_{2}e_{6}=(ap)(xp),&e_{3}e_{8}=(ax)(pz),\\ e_{3}e_{4}=(ax)(bx),&e_{3}e_{7}=(ax)(xq),&e_{3}e_{9}=(ax)qz),&e_{3}e_{6}=(ax)(xp),\\ e_{4}e_{8}=(bx)(pz),&e_{4}e_{7}=(bx)(xq),&e_{4}e_{9}=(bx)(qz),&e_{4}e_{5}=(bx)(bq),\\ e_{4}e_{6}=(bx)(xp),&e_{5}e_{9}=(bq)(qz),&e_{5}e_{6}=(bq)(xp),&e_{5}e_{8}=(bq)(pz),\\ e_{5}e_{7}=(bq)(xq),&e_{6}e_{9}=(xp)(qz),&e_{6}e_{8}=(xp)(pz),&e_{6}e_{7}=(xp)(xq),\\ e_{7}e_{9}=(xq)(qz),&e_{8}e_{9}=(pz)(qz),&e_{2}^{2}=(ap)^{2},&e_{3}^{2}=(ax)^{2},\\ e_{4}^{2}=(bx)^{2},&e_{5}^{2}=(bq)^{2},&e_{6}^{2}=(xp)^{2},&e_{7}^{2}=(xq)^{2},\\ e_{8}^{2}=(pz)^{2},&e_{9}^{2}=(qz)^{2}.&&\end{array}$
Assume $N\geq 3$. Suppose that $I(G)^{N-1}$ has linear quotients with the ordering ${\mathcal{N}}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{N-1}$ and $u_{r}=e_{8}^{N-1}$. Let ${\mathcal{N}}^{(1)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{N}$ coming from ${\mathcal{N}}$ as in Conjecture 3.1. As before, use the following notation to refer to elements in ${\mathcal{N}}^{(1)}$
${\mathcal{U}}e_{1}={\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8}),{\mathcal{A}}(e_{9}),$
where $w_{k}\in{\mathcal{A}}(e_{i})$ means $e_{i}\shortmid w_{k}$ and $e_{j}\nshortmid w_{k}$ for each $j<i$. We claim that the colon ideal $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear for each $w_{n}$. If $w_{n}\in{\mathcal{A}}(e_{1})$, the claim follows from the induction hypothesis.
Let $w_{n}=u_{s}e_{2}\in{\mathcal{A}}(e_{2})$ for some $s\leq r$. Note that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{2})=(b)$. If $w_{i}\in{\mathcal{A}}(e_{2})$, it follows from the induction hypothesis that there exists $j<n$ such that $w_{j}:w_{n}$ is a variable dividing $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})$. If $b\nmid u_{s}$, then $w_{i}:w_{n}\in(b)$ since each element in ${\mathcal{A}}(e_{1})$ is divisible by $b$. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear and we are done. Suppose $b\mid u_{s}$. Then either $e_{4}$ or $e_{5}$ divides $u_{s}$ since $e_{1}\nshortmid u_{s}$. Since $e_{2}e_{4}=e_{1}e_{6}$, it is not possible to have $e_{4}\shortmid u_{s}$ So, $e_{5}\shortmid u_{s}$ where $u_{s}=w^{\prime}e_{5}$ for a generator $w^{\prime}\in I(G)^{N-2}$. In this case, it is important to note that $e_{3}\nshortmid u_{s}$ as $e_{3}e_{5}=e_{1}e_{7}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1},w_{k}=w^{\prime}e_{6}e_{1}$ and $w_{l}=w^{\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a),(w_{k}:w_{n})=(x)$ and $(w_{l}:w_{n})=(z)$. This means $(a,b,x,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
Set $w_{i}=e_{1}^{t_{1}}\cdots e_{9}^{t_{9}}$ where $t_{1}>0$ and $w_{n}=e_{2}^{t^{\prime}_{2}}e_{5}^{t^{\prime}_{5}}e_{6}^{t^{\prime}_{6}}e_{7}^{t^{\prime}_{7}}e_{8}^{t^{\prime}_{8}}e_{9}^{t^{\prime}_{9}}$. If $w_{i}:w_{n}\notin(a,b,x,z)$, we obtain the following inequalities by examining the degrees of these four variables in $w_{i}$ and $w_{n}$
$t_{1}+t_{2}+t_{3}$ $\leq t^{\prime}_{2}$
$t_{1}+t_{4}+t_{5}$ $\leq t^{\prime}_{5}$
$t_{3}+t_{4}+t_{6}+t_{7}$ $\leq t^{\prime}_{6}+t^{\prime}_{7}$
$t_{8}+t_{9}$ $\leq t^{\prime}_{8}+t^{\prime}_{9}$
which implies that $t_{1}+t_{3}+t_{4}\leq 0$, a contradiction. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear for each $w_{n}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$.
Let $w_{n}=u_{s}e_{3}\in{\mathcal{A}}(e_{3})$ for some $s\leq r$. First recall that $e_{1},e_{2}\nshortmid u_{s}$ and also note that $e_{5}\nshortmid u_{s}$ since $e_{3}e_{5}=e_{1}e_{7}$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous case. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup({\mathcal{A}}(e_{2})$. Then, $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:w_{n})=(b)$ and $(u_{s}e_{2}:w_{n})=(p)$. If $b\nmid u_{s}$ and $p\nmid u_{s}$, then $(w_{i}:w_{n})\subseteq(b,p)$ implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. So, the claim holds for this case. If $b\mid u_{s}$, then $e_{4}\shortmid u_{s}$. Let $u_{s}=w^{\prime}e_{4}$ for a generator $w^{\prime}\in I(G)^{q-2}$. Consider $w_{j}=w^{\prime}e_{3}e_{1}$ and $w_{k}=e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ such that $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(q)$. This means $(a,b,p,q)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$. One can verify that the same containment holds for the remaining case $p\mid u_{s}$ (which is split into two cases: $e_{6}\shortmid u_{s}$ or $e_{8}\shortmid u_{s}$).
As in the previous case, let $w_{i}=e_{1}^{t_{1}}\cdots e_{9}^{t_{9}}$ where either $t_{1}>0$ or $t_{2}>0$ and $w_{n}=e_{3}^{t^{\prime}_{3}}e_{4}^{t^{\prime}_{4}}e_{6}^{t^{\prime}_{6}}e_{7}^{t^{\prime}_{7}}e_{8}^{t^{\prime}_{8}}e_{9}^{t^{\prime}_{9}}$. If $w_{i}:w_{n}\notin(a,b,p,q)$, as before, considering degrees of these variables results with $t_{1}+t_{2}+t_{5}\leq 0$, a contradiction. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case as well.
Let $w_{n}=u_{s}e_{4}\in{\mathcal{A}}(e_{4})$ for some $s\leq r$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{4})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{4})$ follows from the induction hypothesis. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that this collection contains all generators of $I(G)^{q}$ that are divisible by $a$ and also $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
Let $w_{n}=u_{s}e_{5}\in{\mathcal{A}}(e_{5})$ for some $s\leq r$. Note that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{5})=(a)$ and $(u_{s}e_{4}:u_{s}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{4})$. Note that each such $w_{i}$ is divisible by $a$ or $x$ and also $a\nmid w_{n}$. If $x\nmid w_{n}$, then $w_{i}:w_{n}\in(a,x)$ and the claim holds. If $x\mid w_{n}$, then either $e_{6}\shortmid u_{s}$ or $e_{7}\shortmid u_{s}$. We consider the $e_{6}\shortmid u_{s}$ case but the other one leads to the same conclusion. Let $u_{s}=w^{\prime}e_{6}$ for a generator $w^{\prime}\in I(G)^{q-2}$. Consider the elements $w_{j}=w^{\prime}e_{4}e_{2}$ and $w_{k}=w^{\prime}e_{6}e_{4}$ where $j,k<n$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(x)$. This means $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear and the claim holds in this case.
[…]
Let $w_{n}=u_{s}e_{7}\in{\mathcal{A}}(e_{7})$ for some $s\leq r$. Notice that $(a,b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{3}:u_{s}e_{7})=(a),(u_{s}e_{4}:u_{s}e_{7})=(b)$ and $(u_{s}e_{6}:u_{s}e_{7})=(p)$. The case $w_{i}\in{\mathcal{A}}(e_{7})$ follows from the induction hypothesis. For the remaining case, note that $w_{i}$ is divisible by either $a$ or $b$ or $p$ while $w_{n}$ is not divisible by neither $a$ or $b$. Furthermore, $w_{n}$ is not divisible by $p$ because if it does, then $e_{8}\shortmid u_{s}$ which implies that $e_{7}e_{8}=e_{6}e_{9}$ divides $w_{n}$, a contradiction. Thus, $w_{i}:w_{n}\in(a,b,p)$ and $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. Hence, the claim holds for this case.
Let $w_{n}=u_{s}e_{8}\in{\mathcal{A}}(e_{8})$ for some $s\leq r$. Notice that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{2}:u_{s}e_{8})=(a)$ and $(u_{s}e_{6}:u_{s}e_{8})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{8})$ follows from the induction hypothesis. For the remaining case, notice that each $w_{i}$ is divisible by either $a,b$ or $x$ while $w_{n}$ is not divisible by any of these three variables. Thus, $w_{i}:w_{n}\in(a,x)$ which means $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
For the final case, we have $w_{n}=e_{9}^{N}$. Observe that $(e_{9}^{N-1}e_{8}:e_{9}^{N})=(p)$ and $(e_{9}^{N-1}e_{5}:e_{9}^{N})=(p)$ and $(e_{9}^{N-1}e_{7}:e_{9}^{N})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{9}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$.
Now, we introduce a peculiar concept of finite graphs.
Definition 4.4.
A finite graph $\Gamma$ is called CDCC graph if $\Gamma$ satisfies the following conditions:
  • $\Gamma$ is gapfree;
  • $\Gamma$ contains (i) cricket, (ii) diamond, (iii) $\mathbf{C}_{4}$ and (iv) $\mathbf{C}_{5}$ as induced subgraphs.
Let $\Gamma$ be a CDCC graph. Since $\Gamma$ contains $C_{5}$, it follows that the edge ideal $I(\Gamma)$ does not have linear resolution. Furthermore, since $\Gamma$ contains cricket, diamond and $C_{4}$, it is unclear if all powers $I(\Gamma)^{t}$ with $t\geq 2$ have linear resolution.
$z$$z^{\prime}$$p$$q$$x$$a$$b$
Figure 4. A CDCC graph with minimum number of vertices
Example 4.5.
It is not difficult to see that no graph on $6$ vertices can satisfy all of the conditions in Definition 4.4. The graph $\Gamma$ displayed in Figure 4 is obtained from the graph in Figure 3 by duplicating the vertex $z$. It is an example of a CDCC graph with smallest number of vertices.
We now come to the first main theorem in the present paper.
Theorem 4.6.
Given $n\geq 7$, there exists a CDCC graph $\Gamma_{n}$ with $n$ vertices for which all powers $I(\Gamma_{n})^{t}$ with $t\geq 2$ have linear quotients.
Proof.
Let $\Gamma_{7}$ denote the CDCC graph of Figure 4. Since a finite graph obtained by duplicating a vertex from a CDCC graph is again CDCC graph, the desired result follows from Propositions 2.2 and 4.3.
While the matching number of the graph in Figure 4 is $3$; in general, the matching number of a CDCC graph can be arbitrarily large. Indeed, duplicating both endpoints of an edge increases the matching number without changing the existing induced subgraphs and gapfreeness.
5. Vertex duplication and linear quotients of powers
The techniques in the prior sections can be expanded to produce additional classes of graphs whose powers have linear quotients. In Section 2 it was shown that linear quotients are preserved under vertex duplication. There is a second way to define vertex duplication, where the vertex and its duplicate are connected via an edge. In this section, we show that for a gapfree graph, this type of vertex duplication preserves linear quotients as long as the duplication preserves gapfree. We first fix notation for the two types of duplication.
Notation 5.1.
Let $G$ be a gap free graph. Fix a vertex $x\in V(G)$. Define $G^{x}$ to be the graph obtained from $G$ by duplicating $x$ as in Definition2.1. Let
$V(G^{x})=V(G)\cup\{y\}$
for some $y\not\in V(G)$. Define $H^{x}$ to be the graph with vertex set $V(H^{x})=V(G^{x})$ and $E(H)=E(G^{x})\cup\{\{x,y\}\}$.
For any $W\subseteq V(H^{x})=V$ and any monomial $w\in I(H^{x})^{s}$, define $w_{W}$ to be the monomial formed by localizing $w$ at $W$. That is, if $w=\prod_{v_{i}\in V}v_{i}^{s_{i}}$, then $w_{W}=\prod_{v_{i}\in W}v_{i}^{s_{i}}$.
Next, we state some useful facts regarding the duplicated vertices. The proofs are straightforward from the definitions and so are not included.
Lemma 5.2.
Let $G$ be a gapfree graph. Using 5.1,
  • $N_{G^{x}}(x)=N_{G^{x}}(y)$.
  • $N_{H^{x}}[x]=N_{H^{x}}[y]$
  • $H^{x}$ is gap-free if and only if $V(H^{x})\setminus N_{H^{x}}[x]$ is an independent set.
We are now ready to prove the second main theorem in the present paper.
Theorem 5.3.
Let $G$ be a gapfree graph and fix $x\in V(G)$. Using 5.1, assume $H^{x}$ is gapfree and and $I(G)^{s}$ has linear quotients for some $s\geq 1$. Then $I(H^{x})^{s}$ has linear quotients.
Proof.
For any generator $w$ of $I(H^{x})^{s}$ define
$\mu(w)=\min\{i:\frac{w}{(xy)^{i}}\in Gens(I(G^{x})^{s-i})\}.$
Note that $\mu(w)=0$ if and only if $w$ is a minimal generator of $I(G^{x})^{s}$. Set $V=V(H^{x})$. Partition $V$ into sets:
$Z$ $=$ $\{x,y\}$
$A$ $=$ $N_{G}(x)$
$B$ $=$ $V(G)\setminus N_{G}[x]$
By 5.2, $B$ is an independent set. Fix an order $>$ on the elements of $B$. By Proposition 2.2 there is a linear quotients ordering $w_{1},w_{2},\dots,w_{q}$ on the minimal generators of $I(G^{x})^{s}$. We will extend this order to a linear quotients ordering $w_{1},\dots,w_{q},w_{q+1},\dots,w_{r}$ on the minimal generators of $I(H^{x})^{s}$. To simplify the notation, we adopt the convention that $w_{1}>w_{2}>\dots>w_{r}$. For generators $w$ and $w^{\prime}$ of $I(H^{x})^{s}$ not belonging to $I(G^{x})^{s}$, set $w>w^{\prime}$ if:
  1. (1)
    $\mu(w)<\mu(w^{\prime})$, or
  2. (2)
    $\mu(w)=\mu(w^{\prime})$ and $\deg(w_{Z})<\deg(w^{\prime}_{Z})$, or
  3. (3)
    $\mu(w)=\mu(w^{\prime})$, $\deg(w_{Z})=\deg(w^{\prime}_{Z})$, and $|\deg_{x}(w)-\deg_{y}(w)|<|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|$, or
  4. (4)
    $\mu(w)=\mu(w^{\prime})$, $\deg(w_{Z})=\deg(w^{\prime}_{Z})$, $|\deg_{x}(w)-\deg_{y}(w)|=|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|$, and $w_{B}>w^{\prime}_{B}$ in the lexicographic ordering on $B$, using the order fixed above.
Let $w=w_{i}$ for some $i\geq 2$. We will show that $J=(w_{1},\dots,w_{i-1}):w_{i}$ is generated by variables. We may assume that $\mu(w)\neq 0$ by definition of the order and Proposition 2.2. Since $B$ is an independent set, the edges of $H^{x}$ can be partitioned as follows:
$E_{1}$ $=$ $\{xy\}$
$E_{2}$ $=$ $\{ab\mid a\in A,b\in B\}$
$E_{3}$ $=$ $\{az\mid a\in A,z\in Z\}$
$E_{4}$ $=$ $\{aa^{\prime}\mid a,a^{\prime}\in A\}.$
We make two key observations which will be used in the proofs.
  • (i)
    Let $w$ be a generator of $I(H^{x})^{s}$ with $\mu(w)\neq 0$. For any representation
    $w=(xy)^{\mu(w)}e_{1}\dots e_{s-\mu(w)}$
    for edges $e_{1},\dots,e_{s-\mu(w)}$, we have $e_{j}\notin E_{4}$ for all $j=1,\dots,s-\mu(w)$.
  • (ii)
    $A$ is a subset of the set of generators of $J$.
Proof of (i). Assume to the contrary that $e_{1}=a_{1}a_{2}$ for some $a_{1},a_{2}\in A$. Then we can write
$w=(xy)^{\mu(w)-1}(xa_{1})(ya_{2})e_{2}\dots e_{s-\mu(w)}$
which is a contradiction to the minimality of $\mu(w)$.
Proof of (ii). Let $a\in A$. Then $u:=\frac{w}{xy}(xa)$ is a minimal generator of $I(H^{x})^{s}$ with $\mu(u)\leq\mu(w)-1$. Therefore $u>w$ and $u:w=a\in J$, proving the claim.
Let $v>w$. We will show that there is a variable generator of $J$ dividing $v:w$.
CASE 1: Suppose that $\mu(v)<\mu(w)$. Using observation (i), each $e_{i}$ in the representation of $w$ either has precisely one endpoint in $A$ or is in $E_{1}$. It follows that
$\deg(v_{A})\geq s-\mu(v)>s-\mu(w)=\deg(w_{A}).$
Therefore, there is an $a\in A$ such that $a$ divides $v_{A}:w_{A}$. Since $v_{A}:w_{A}$ divides $v:w$, the proof follows from observation (ii).
CASE 2: Suppose that $1\leq\mu(v)=\mu(w)$. As above, from observation (i) it follows that
$\deg(v_{A})=s-\mu(v)=s-\mu(w)=\deg(w_{A}).$
Now by observation (ii), we may assume for every $a\in A$ we have $\deg_{a}(v)\leq\deg_{a}(w)$. However, since
$\deg(v_{A})=\sum_{a\in A}\deg_{a}(v)\leq\sum_{a\in A}\deg_{a}(w)=\deg(w_{A})$
the inequality cannot be strict for any $a\in A$. Thus for the remainder of the proof, we assume that $v_{A}=w_{A}$.
Case 2.1: Suppose that $\deg(v_{Z})<\deg(w_{Z})$. Let $\gamma=\mu(v)=\mu(w)$. By observation (i) we may represent $v=(xy)^{\gamma}e_{1}\dots e_{m}f_{1}\dots f_{k}$ and $w=(xy)^{\gamma}g_{1}\dots g_{m^{\prime}}h_{1}\dots h_{k^{\prime}}$ with $e_{j},g_{j}\in E_{3}$ and $f_{j},h_{j}\in E_{2}$ for all $j$. Note that
$2\gamma+m=\deg(v_{Z})<\deg(w_{Z})=2\gamma+m^{\prime}.$
On the other hand, since $m+k=m^{\prime}+k^{\prime}$, it follows that $k>k^{\prime}$. Therefore
$\deg((f_{1}\dots f_{k})_{A})=k>k^{\prime}=\deg((h_{1}\dots h_{k^{\prime}})_{A}).$
Since $v_{A}=w_{A}$ and from the above inequality, there exists $a\mid(f_{1}\dots f_{k})_{A}$ such that $a\mid(g_{1}\dots g_{m^{\prime}})_{A}$. Without loss of generality, we may assume that $f_{1}=ab$ for some $b\in B$ and $g_{1}=az$ for some $z\in Z$. If $b\mid(v:w)$, set $w^{\prime}:=\frac{w}{g_{1}}f_{1}=\frac{w}{z}b$. Then $w^{\prime}>w$ with $w^{\prime}:w=b$, which completes the proof.
Assume that $b\nmid(v:w)$. Then, without loss of generality, we may assume that $h_{1}=a^{\prime}b$ for some $a^{\prime}\in A$. Then we can re-write $w$ as follows:
$w=(xy)^{\gamma}(a^{\prime}z)g_{2}\dots g_{m^{\prime}}(ab)h_{2}\dots h_{k^{\prime}}.$
Then, the same argument can be applied recursively to $v/(ab)$ and $w/(ab)$.
Case 2.2: Suppose that $\deg(v_{Z})=\deg(w_{Z})$.
Case 2.2-(i): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|<|\deg_{x}(w)-\deg_{y}(w)|$. Since the right side of the inequality cannot be 00, without loss of generality, assume that $\deg_{x}(w)>\deg_{y}(w)$. Then since $\deg(v_{Z})=\deg(w_{Z})$, we have $\deg_{y}(v)>\deg_{y}(w)$ and $y\mid v:w$. Since $\deg_{x}(w)>\deg_{y}(w)$ there is an $a\in A$ with $e_{1}=ax$ in the representation of $w$. By definition of $H^{x}$, $a\in N(y)$ also. Set $w^{\prime}:=\frac{w}{ax}ay$. Then $w^{\prime}>w$ since $|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|<|\deg_{x}(w)-\deg_{y}(w)|$, and $w^{\prime}:w=y$.
Case 2.2-(ii): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|=|\deg_{x}(w)-\deg_{y}(w)|$. Assume for a moment that $v_{Z}\neq w_{Z}$. Then without loss of generality, we may write
  • $\deg_{x}(v)=r_{1}=\deg_{y}(w)$
  • $\deg_{x}(w)=r_{2}=\deg_{y}(v)$
for some integers $r_{1}>r_{2}$. Then $w^{\prime}:=\frac{w}{y}x>w$ and $x=w^{\prime}:w$ divides $v:w$. Therefore, we may assume $v_{Z}=w_{Z}$. For the sake of clarity, we remind that the following statements are assumed at this point of the proof:
  • $1\leq\mu(v)=\mu(w)=\gamma$
  • $v_{A}=w_{A}$
  • $v_{Z}=w_{Z}$.
We fix a presentation $v=(xy)^{\gamma}f_{1}f_{2}\cdots f_{t}$ and $w=(xy)^{\gamma}g_{1}g_{2}\cdots g_{t}$. For simplicity, we assume that $f_{i_{1}}\dots f_{i_{m}}\neq g_{j_{1}}\dots g_{j_{m}}$ for all $m$-fold products with $1\leq m\leq t$ since otherwise the argument below is still valid recursively after removing the common factors.
Fix $p$ greatest in the ordering fixed on $B$ such that $p\mid(v:w).$ We claim that there is $u>w$ such that $u:w=p$. After rearranging the indices, we may write $f_{1}=a_{1}p$ for some $a_{1}\in A$.
Step 1: If there exists $\delta_{1}\in Z$ or $\delta_{1}\in B$ with $p>\delta_{1}$ such that $g_{1}=a_{1}\delta_{1}$ after some rearrangement of indices, set
$u=\frac{w}{g_{1}}f_{1}=\frac{w}{(a_{1}\delta_{1})}(a_{1}p)$
Then $u>w$ in the ordering since $\deg(u_{Z})<\deg(w_{Z})$ in the first case and $u_{B}>w_{B}$ in the latter one. Therefore, we may assume that $g_{1}=a_{1}\delta_{1}$ for some $\delta_{1}\in B$ with $\delta_{1}>p$.
Step 2: By the choice of $p$, we may set $f_{2}=a_{2}\delta_{1}$ for some $a_{2}\in A$. As before, if, up to reordering, $g_{2}=a_{2}\delta_{2}$ for some $\delta_{2}\in Z$ or $\delta_{2}\in B$ with $p>\delta_{2}$ we form the monomial
$u=\frac{w}{g_{1}g_{2}}f_{1}f_{2}=\frac{w}{(a_{1}\delta_{1})(a_{2}\delta_{2})}(a_{1}p)(a_{2}\delta_{1}).$
which satisfies the desired property. Therefore, we may assume that $\delta_{2}\in B$ and $\delta_{2}>p$.
Step 3: By the choice of $p$, we may set $f_{3}=a_{3}\delta_{2}$ for some $a_{3}\in A$. As before, if, up to reordering, $g_{3}=a_{3}\delta_{3}$ for some $\delta_{3}\in Z$ or $\delta_{3}\in B$ with $p>\delta_{3}$ we form the monomial
$u=\frac{w}{g_{1}g_{2}g_{3}}f_{1}f_{2}f_{3}=\frac{w}{(a_{1}\delta_{1})(a_{2}\delta_{2})(a_{3}\delta_{3})}(a_{1}p)(a_{2}\delta_{1})(a_{3}\delta_{2}).$
which satisfies the desired property. Therefore, we may assume $\delta_{3}\in B$ and $\delta_{3}>p$. The process must stop after $k$ steps when there is no $\delta_{k}\in B$ left with $\delta_{k}>p$.
Corollary 5.4.
Suppose that $I(G)^{s}$ has linear quotients for some $s$. Let $x$ be a vertex of $G$ such that non-neighbours of $x$ is an independent set. Let $H$ be the graph obtained from $G$ by replacing $x$ with a clique. Then $I(H)^{s}$ has linear quotients.
As an example of how this result can be applied, consider the class of graphs formed by replacing the central vertex $x$ of Figure 2 by a complete graph $K_{n}$, with each vertex of $K_{n}$ connected to $a,b,z$. We will denote such a graph by $(C_{5},K_{n})$ and will label the vertices of $K_{n}$ by $x_{1},\ldots,x_{n}$. The case $n=4$ is shown below.
$z$$p$$q$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$a$$b$
Figure 5. $(C_{5},K_{4})$
Example 5.5.
Let $I$ be the edge ideal of the graph $(C_{5},K_{n})$ for some $n\geq 1$. Then $I^{s}$ has linear quotients for all $s\geq 2$. The case $n=1$ holds by 4.2. For $n>1$ the result holds via induction using Proposition 2.2.
Notice that the above argument cannot extend to replace any vertex of the exterior pentagon with an edge. In particular, this would result in a finite graph that is not gapfree. However, the technique can be applied to the central vertex in Figure 3, for instance.
6. A partial answer to Conjecture 3.1
In the present section, we are faced with a challenge on a slightly modified version of Conjecture 3.1.
Definition 6.1.
Let $G$ be a finite graph on the vertex set $V(G)$ and $E(G)$ the set of its edges. An admissible edge ordering of $G$ is a total order $\succ$ of the edges in $G$ with following property: for any $\{a,b\},\{c,d\}\in E(G)$ such that $\{a,b\}\cap\{c,d\}=\emptyset$ and $\{a,b\}\succ\{c,d\}$, we have either all edges incident to $a$ are larger than $\{c,d\}$ or all edges incident to $b$ are larger than $\{c,d\}$.
In what follows, unless there is a misunderstanding, we use the notation $ab$ instead of $\{a,b\}$ for an edge of a finite graph.
Lemma 6.2.
Every finite graph possesses an admissible ordering of the edges.
Proof.
We shall use induction on $n=|V(G)|$, the number of vertices in $G$. The assertion is trivial if $n=1$ or $n=2$. Suppose that $n\geq 3$ and the assertion is known to hold for any graph with fewer than $n$ vertices. Let $x\in V(G)$ be any vertex in $G$, and set $G^{\prime}=G\setminus\{x\}$. Suppose that the edges incident to $x$ in $G$ are $xx_{1},\dots,xx_{s}$.
By the induction hypothesis, there exists an admissible ordering of the edges in $G^{\prime}$. We shall order the edges in $G$ by $xx_{1}\succ\dots\succ xx_{s}\succ\text{ edges in }G^{\prime}$. We will show that this is an admissible ordering of the edges in $G$. Consider any two edges $ab>cd$ with $ab\cap cd=\emptyset$ in this ordering. Clearly, $cd$ cannot be incident to $x$.
Consider the case that $ab$ is not incident to $x$. Then, both $ab$ and $cd$ are edges in $G^{\prime}$. Thus, we can suppose that all edges incident to $a$ in $G^{\prime}$ are larger than $cd$. If $a\not\in\{x_{1},\dots,x_{s}\}$, then edges incident to $a$ in $G$ and $G^{\prime}$ are the same, and we have the required property. If $a\in\{x_{1},\dots,x_{s}\}$, then the only additional edge incident to $a$ in $G$ would be $xa$, which is larger than $cd$, by the edge ordering.
Consider the remaining case that $ab$ is incident to $x$. Without loss of generality, we may assume that $a=x$. In this case, it is clear that all edges incident to $a$ are larger than $cd$.
From now on, for a finite graph $G$, fix an admissible order of its edges:
($\ast$) $e_{1}\succ\dots\succ e_{s}$
The existence of this ordering is due to 6.2.
Definition 6.3.
Let $G$ be a finite graph with edge ideal $I=I(G)$. Let
($\diamondsuit$) $g_{1}>\cdots>g_{k}$
be an order of the generators of $I^{2}$. We recursively construct the following orders for the generators of powers of $I$: if, for $q\geq 2$,
${\mathcal{M}}^{(q)}:u_{1}>\dots>u_{r}$
denotes the constructed (or known) order of the generators of $I(G)^{q}$, then the order of the generators of $I(G)^{q+1}$ is given by
${\mathcal{M}}^{(q+1)}:u_{1}e_{1}>\dots>u_{r}e_{1}>u_{1}e_{2}>\dots>u_{r}e_{2}>\dots>u_{r}e_{s},$
where a generator $u_{i}e_{j}$ is omitted (redundant) if it has already appeared before in the ordering. We call these orders ${\mathcal{M}}^{(q)}$ of the generators of $I^{q}$ for $q=3,4,5,\ldots$ are the ($\diamondsuit$6.3)–($\ast$6) compatible orders.
We now come to the third main theorem in the present paper.
Theorem 6.4.
Let $G$ be a finite graph on the vertex set $V(G)$ and $E(G)$ the set of its edges. Let $I=I(G)$ be the edge ideal of $G$. Suppose that $I^{2}$ has linear quotients with an order of the generators written in the form of (italic-$\diamondsuit$6.3). Suppose also that, for some integer $q\geq 7$, $I^{2},\dots,I^{q}$ have linear quotients with the (italic-$\diamondsuit$6.3)–(italic-$\ast$6) compatible orders of their generators. Then, $I^{q+1}$ also has linear quotients with the (italic-$\diamondsuit$6.3)–(italic-$\ast$6) compatible order of its generators.
Proof.
For simplicity of the notation, for monomials $M$ and $N$, we shall write $M:N$ for $M/\operatorname{lcm}(M,N)$, which is the monomial generator of the colon ideal $(M):(N)$. Consider any two generators $A=u_{i}e_{j}>u_{k}e_{\ell}=B$ of $I^{q+1}$ in ${\mathcal{M}}^{(q+1)}$, and assume that they appear in the ordering ${\mathcal{M}}^{(q+1)}$ the first time in this representation; that is, $e_{j}$ and $e_{\ell}$ are the largest among all edges in the decomposition of $A$ and $B$, respectively. Set
$Q=A:B.$
We need to show that there exists a generator $C$ of $I^{q+1}$ such that $C>B$ in ${\mathcal{M}}^{(q+1)}$, and $C:B$ is a variable that divides $Q$.
Since $A>B$ in ${\mathcal{M}}^{(q+1)}$, we can assume that $e_{j}\succeq e_{\ell}$; that is, $j\leq\ell$. If $j=\ell$, then $u_{i}>u_{k}$ in ${\mathcal{M}}^{(q)}$, so there exists $u_{t}>u_{k}$ in $I^{q}$ such that $u_{t}:u_{k}$ is a variable that divides $Q=u_{i}:u_{k}$. Let $C=u_{t}e_{\ell}$. Then, $C:u_{k}e_{\ell}=u_{t}:u_{k}$ is a variable that divides $Q$. Since $B$ appears in ${\mathcal{M}}^{(q+1)}$ first as $u_{k}e_{\ell}$ and $u_{i}>u_{k}$ in ${\mathcal{M}}^{(q)}$, it follows that $C>B$ in ${\mathcal{M}}^{(q+1)}$ and the assertion is proved in this case.
We can assume that $e_{i}\not=e_{\ell}$, i.e., $j\not=\ell$. Write $u_{i}=L_{1}\cdots L_{q}$ and $u_{k}=M_{1}\cdots M_{q}$, where $L_{i}$’s and $M_{j}$’s are edges in $G$, and set $e_{i}=ab,e_{\ell}=cd$. As observed before, $ab\succeq L_{t}$ and $cd\succeq M_{t}$ for all $t=1,\dots,q$. Also, $ab\succ cd$. We shall analyze the following two possibilities: $ab\cap cd=\emptyset$ and $ab\cap cd\not=\emptyset$.
Case 1: $ab\cap cd=\emptyset$. It can be seen that $ab\nmid u_{k}$. Indeed, if $ab\mid u_{k}$ then there must exist edges $ae=M_{t^{\prime}}$ and $bf=M_{t^{\prime\prime}}$ in $u_{k}$, for some $t^{\prime},t^{\prime\prime}$ (it is possible that $M_{t^{\prime}}=M_{t^{\prime\prime}}=ab$). However, by the property of an admissible ordering, we must have either $ae\succ cd$ or $bf\succ cd$, which is a contradiction to the assumption that $M_{t}\preceq cd$ for all $t=1,\dots,q$.
If $a\mid u_{k}$ and $b\nmid u_{k}$ (similarly for the case where $a\nmid u_{k}$ and $b\mid u_{k}$), then there exists an edge $ae$ in the presentation of $u_{k}$. Without loss of generality, we may assume that $ae=M_{q}$. Now, set
$C=(M_{1}\cdots M_{q-1}\cdot cd)\cdot ab.$
It can be seen that, since $ab\succ cd$, $C>B$ in ${\mathcal{M}}^{(q+1)}$. Moreover, $C:B=b$, which divides $Q$, as desired.
If $a\nmid u_{k}$ and $b\nmid u_{k}$ then set
$u_{l}=M_{1}\cdots M_{q-1}\cdot ab.$
Since $ab\succ cd\succeq M_{q}$, $u_{l}>u_{k}$ in ${\mathcal{M}}^{(q)}$. Therefore, by the linear quotients of $I^{q}$, there exists a generator $u_{t}>u_{k}$ of $I^{q}$ such that $u_{t}:u_{k}$ is a variable that divides $u_{l}:u_{k}$. Since $u_{l}:u_{k}=ab$, we have $u_{t}:u_{k}$ is either $a$ or $b$. Set
$C=u_{t}\cdot cd.$
Since $B$ appears first in ${\mathcal{M}}^{(q+1)}$ in the form $u_{k}\cdot cd$, we have $C>B$. Moreover, $C:B$ is either $a$ or $b$, which divides $Q$ as desired.
Case 2: $ab\cap cd\not=\emptyset$. Since $ab\not=cd$, we may assume that $a=c$ and $b\not=d$.
If $b\mid Q$, then set
$C=u_{k}\cdot ab.$
Clearly, $C>B$ since $ab\succ ad$, and $C:B=b$ which divides $Q$.
If $a\mid Q$ and $b\nmid Q$, then there must exist an edge $be$ in $u_{k}$. Without loss of generality, assume that $be=M_{q}$ (particularly, $e\not=a$). Set
$C=(M_{1}\cdots M_{q-1}\cdot ad)\cdot ab.$
Since $ab\succ ad\succeq be$, we have $C>B$. Also, $C:B=a$, which divides $Q$.
It remains to consider the case where $a\nmid Q$ and $b\nmid Q$. As before, since $b\nmid Q$, there must exists an edge $be$ in $u_{k}$, and we can assume that $be=M_{q}$.
Suppose that there exists an edge $uv\in E$ such that $uv\mid Q$. Set
$u_{g}=(M_{1}\cdots M_{q-2})\cdot uv\cdot ab\,\,\,\text{ and }\,\,\,u_{h}=(M_{1}\cdots M_{q-2})\cdot be\cdot ad.$
Again, since $ab\succ ad\succeq M_{t}$ for all $t$, we have $u_{g}>u_{h}\geq u_{k}$ in ${\mathcal{M}}^{(q)}$. By the linear quotients of $I^{q}$, there exists a generator $u_{t}>u_{h}$ of $I^{q}$ such that $u_{t}:u_{h}$ is a variable that divides $u_{g}:u_{h}=uv:de$ (which is necessarily not 1). Set
$C=u_{t}\cdot M_{q-1}.$
Since $C$ contains $ab$, we have $C>u_{h}\cdot M_{q-1}=u_{k}\cdot ad=B$. Furthermore, $C:B=u_{t}:u_{h}$ is a variable which divides $uv\mid Q$. Hence, we can assume now that such an edge $uv$ does not exist. Particularly, it follows that $Q$ is the product of variables (with possibly repetitions) in an independent set of $G$.
Write $Q=x_{1}\dots x_{\gamma}$, where the $x_{l}$’s (with possible repetition) form an independent set in $G$. By relabeling, if necessary, we may assume that $L_{l}=x_{l}y_{l}$, for $l=1,\dots,\gamma$ (in particular, $\gamma\leq q$). Since $y_{l}\nmid Q$, for $l=1,\dots,\gamma$, we may also assume that $M_{l}=y_{l}z_{l}$, for $l=1,\dots,\gamma$, or that some of the $y_{l}$’s are $d$ or $e$.
If $x_{l}d\in E$ (similarly, if $x_{l}e\in E$) for some $l\in\{1,\dots,\gamma\}$, then set
$C=(M_{1}\cdots M_{q-1})\cdot x_{l}d\cdot ab.$
As observed before, since $ab\succ ad\succeq M_{t}$ for all $t$, $C>B$. Also, $C:B=x_{l}$, which divides $Q$. Suppose that $x_{l}d,x_{l}e\not\in E$ for any $l=1,\dots,\gamma$. This, by considering $\deg A=\deg B$, implies that $\gamma<q$.
Case 2a. Consider the case where there exists $\theta\in\{\gamma+1,\dots,q\}$ such that $L_{\theta}\subseteq Z=\{z_{1},\dots,z_{\gamma},d,e\}$.
If $L_{\theta}=de$, then $B=(M_{1}\cdots M_{q-1})\cdot L_{\theta}\cdot ab$, which already appeared before in ${\mathcal{M}}^{(q+1)}$, a contradiction. If, for instance, $L_{\theta}=z_{1}d$ (similarly for the case where $L_{\theta}=z_{1}e$), then set
$C=(M_{2}\cdots M_{q-1})\cdot L_{1}\cdot L_{\theta}\cdot ab.$
Again, we have $C>B$. Also, $C:B=x_{1}:e$, so either $B=C$ is redundant or $C:B=x_{1}$ divides $Q$.
If, for example, $L_{\theta}=\{z_{1},z_{2}\}$, then without loss of generality assume that $M_{q-1}$ is the largest among $M_{3},\dots,M_{q-1}$ and set
$u_{g}=(M_{3}\cdots M_{q-2})\cdot L_{1}\cdot L_{2}\cdot L_{\theta}\cdot ab.$
Clearly, $u_{g}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$. Furthermore, $u_{g}:B/M_{q-1}=x_{1}x_{2}:de$ (which is necessarily not 1). It follows that either $u_{g}=B/M_{q-1}$, which then implies that $B=u_{g}M_{q-1}$ is redundant, or there exists a generator $u_{h}>B/M_{q-1}$ of $I^{q}$ ($u_{h}$ could be $u_{g}$) such that $u_{h}:B/M_{q-1}$ is either $x_{1}$ or $x_{2}$, which then implies that $u_{h}M_{q-1}:B$ is either $x_{1}$ or $x_{2}$ that divides $Q$. This case is completed by showing that $u_{h}M_{q-1}>B$, which is done in the following Claim.
Claim. $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$.
Proof of Claim. Since $u_{h}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$, either $u_{h}$ contains an edge $f$ of $G$ such that $f\succ ad$ or $u_{h}=u_{h}^{\prime}\cdot ad$, where $u_{h}^{\prime}$ is a generator of $I^{q-1}$ and $u_{h}^{\prime}>B/(M_{q-1}\cdot ad)=M_{1}\cdot M_{q-2}\cdot M_{q}$ in ${\mathcal{M}}^{(q-1)}$. If $u_{h}$ contains $f\succ ad$, then $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$ as claimed.
Suppose that $u_{h}=u_{h}^{\prime}\cdot ad$ and $u_{h}^{\prime}>M_{1}\cdots M_{q-2}\cdot M_{q}=N$ in ${\mathcal{M}}^{(q-1)}$. To proceed, without loss of generality, we may assume that the relative orders among the edges $M_{1},M_{2}$ and $M_{q}$ is $M_{1}\succeq M_{2}\succeq M_{q}$.
If $M_{q-1}\succeq M_{1}$ then $u_{h}^{\prime}M_{q-1}>N\cdot M_{q-1}$ in ${\mathcal{M}}^{(q)}$, and so $u_{h}M_{q-1}=u_{h}^{\prime}M_{q-1}\cdot ad>N\cdot M_{q-1}\cdot ad=B$ in ${\mathcal{M}}^{(q+1)}$ as asserted.
If $M_{1}\succ M_{q-1}\succeq M_{2}$, then since $u_{h}^{\prime}>N$ in ${\mathcal{M}}^{(q-1)}$, either $u_{h}^{\prime}$ contains an edge $f^{\prime}\succ M_{1}$ or $u_{h}^{\prime}=M_{1}\cdot u_{h}^{\prime\prime}$, where $u_{h}^{\prime\prime}>M_{2}\cdots M_{q-2}\cdot M_{q}=N^{\prime}$ in ${\mathcal{M}}^{(q-2)}$. In the former case, when $u_{h}^{\prime}$ contains an edge $f^{\prime}\succ M_{1}$, we have $u_{h}^{\prime}M_{q-1}>NM_{q-1}$ in ${\mathcal{M}}^{(q)}$, and so $u_{h}M_{q-1}=u_{h}^{\prime}M_{q-1}\cdot ad>NM_{q-1}\cdot ad=B$ in ${\mathcal{M}}^{(q+1)}$. In the latter scenario, when $u_{h}^{\prime}=M_{1}\cdot u_{h}^{\prime\prime}$ with $u_{h}^{\prime\prime}>N^{\prime}$ in ${\mathcal{M}}^{(q-2)}$, we have $u_{h}^{\prime\prime}M_{q-1}>N^{\prime}M_{q-1}$ in ${\mathcal{M}}^{(q-1)}$, and so $u_{h}^{\prime\prime}M_{q-1}M_{1}>N^{\prime}M_{q-1}M_{1}=NM_{q-1}$ in ${\mathcal{M}}^{(q)}$. It follows that $u_{h}M_{q-1}=u_{h}^{\prime}\cdot M_{q-1}\cdot ad=u_{h}^{\prime\prime}\cdot M_{1}\cdot M_{q-1}\cdot ad>NM_{q-1}\cdot ad=B$ in ${\mathcal{M}}^{(q+1)}$ as desired.
If $M_{1}\succeq M_{2}\succ M_{q-1}\succeq M_{q}$, then since $u_{h}^{\prime}>N$ in ${\mathcal{M}}^{(q-1)}$, either $u_{h}^{\prime}$ contains an edge $f^{\prime}\succ M_{1}$ or $u_{h}^{\prime}=M_{1}\cdot u_{h}^{\prime\prime}$ with $u_{h}^{\prime\prime}>N^{\prime}$ in ${\mathcal{M}}^{(q-2)}$. As before, if $u_{h}^{\prime}$ contains such an $f^{\prime}$ then the claim is proved. On the other hand, since $u_{h}^{\prime\prime}>N^{\prime}$ in ${\mathcal{M}}^{(q-2)}$, either $u_{h}^{\prime\prime}$ contains an edge $f^{\prime\prime}\succ M_{2}$ or $u_{h}^{\prime\prime}=M_{2}\cdot u_{h}^{\prime\prime\prime}$ with $u_{h}^{\prime\prime\prime}>M_{3}\cdots M_{q-2}\cdot M_{q}=N^{\prime\prime}$ in ${\mathcal{M}}^{(q-3)}$. If such an edge $f^{\prime\prime}$ in $u_{h}^{\prime\prime}$ exists, then for the same reason as with $f^{\prime}$, we have $u_{h}^{\prime\prime}M_{q-1}>N^{\prime}M_{q-1}$, and so $u_{h}^{\prime}M_{q-1}=u_{h}^{\prime\prime}M_{q-1}M_{1}>N^{\prime}M_{q-1}M_{1}=NM_{q-1}$, which implies that $u_{h}M_{q-1}=u_{h}^{\prime}M_{q-1}\cdot ad>NM_{q-1}\cdot ad=B$, as asserted. If $u_{h}^{\prime\prime}=M_{2}\cdot u_{h}^{\prime\prime\prime}$ with $u_{h}^{\prime\prime\prime}>N^{\prime\prime}$, then $u_{h}^{\prime\prime\prime}M_{q-1}>N^{\prime\prime}M_{q-1}$, and so $u_{h}^{\prime\prime}M_{q-1}=u_{h}^{\prime\prime\prime}M_{q-1}M_{2}>N^{\prime\prime}M_{q-1}M_{2}=N^{\prime}M_{q-1}$. It follows that $u_{h}^{\prime}M_{q-1}=u_{h}^{\prime\prime}M_{q-1}M_{1}>N^{\prime}M_{q-1}M_{1}=NM_{q-1}$ and, therefore, $u_{h}M_{q-1}=u_{h}^{\prime}M_{q-1}\cdot ad>NM_{q-1}\cdot ad=B$ as claimed.
Finally, consider the case that $M_{q}\succ M_{q-1}$. The argument goes similarly as in the previous cases, where we successively write $u_{h}=u_{h}^{\prime}\cdot ad$, $u_{h}^{\prime}=u_{h}^{\prime\prime}\cdot M_{1}$, $u_{h}^{\prime\prime}=u_{h}^{\prime\prime\prime}\cdot M_{2}$ and $u_{h}^{\prime\prime\prime}=u_{h}^{\prime\prime\prime\prime}\cdot M_{q}$ with $u_{h}^{\prime\prime\prime\prime}>M_{3}\cdots M_{q-2}$ in ${\mathcal{M}}^{(q-4)}$; note that $q-4>2$ since $q\geq 7$. The argument then completes by tracing back as seen before. The Claim is established.
Case 2b. Consider the case where there does not exist such $L_{\theta}\subseteq Z$. Since $L_{\gamma+1}\cdots L_{q}$ is contained in $M_{\gamma+1}\cdots M_{q-1}\cdot d\cdot e\cdot z_{1}\cdots z_{\gamma}$, and $\deg(L_{\gamma+1}\cdots L_{q})=\deg(M_{\gamma+1}\cdots M_{q-1})+2$, there must exist $\theta,\delta\in\{\gamma+1,\dots,q\}$ and $\sigma\in\{\gamma+1,\dots,q-1\}$ such that: $L_{\theta}\cap Z\not=\emptyset,L_{\delta}\cap Z\not=\emptyset$, and $M_{\sigma}$ covers (consists of) two vertices, one from each of the edges $L_{\theta}$ and $L_{\delta}$. For simplicity of notations, we shall assume that $\sigma=3$.
Set $L_{\theta}=uv$ and $L_{\delta}=u^{\prime}v^{\prime}$, where, without loss of generality, $u,u^{\prime}\in Z$, $v,v^{\prime}\not\in Z$, and $vv^{\prime}=M_{3}$. We shall examine the following possibilities.
(a) $u,u^{\prime}\in\{z_{1},\dots,z_{\gamma}\}$; for instance, $u=z_{1}$ and $u^{\prime}=z_{2}$. Set
$u_{g}=(L_{1}L_{2}L_{\theta}L_{\delta})\cdot(M_{4}\cdots M_{q-2})\cdot ab.$
Then, $u_{g}>(M_{1}\cdots M_{q-2})\cdot M_{q}\cdot ad=B/M_{q-1}$ are generator of $I^{q}$. Also, $u_{g}:B/M_{q-1}=x_{1}x_{2}:de$ (which is necessarily not 1). Thus, there exists $u_{h}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$ so that $u_{h}:B/M_{q-1}$ is a variable that divides $x_{1}x_{2}$. By a similar argument as in the Claim, in which we consider the relative order of $M_{q-1}$ compared to $M_{1},M_{2},M_{3}$ and $M_{q}$ noting that $q-5\geq 2$ since $q\geq 7$, we see that $u_{h}M_{q-1}>B$. Also, $u_{h}M_{q-1}:B$ is a variable that divides $Q$ as desired.
(b) $u\in\{z_{1},\dots,z_{\gamma}\}$ and $u^{\prime}\in\{d,e\}$; for instance, $u=z_{1}$ and $u^{\prime}=e$. Set
$C=(L_{1}L_{\theta}L_{\delta}M_{2})\cdot(M_{4}\cdot M_{q-1})\cdot(ab).$
Then, $C>B$, and $C:B=x_{1}:d$. Thus, either $B=C$ is redundant, or $C:B=x_{1}$, which divides $Q$.
The next two propositions show that linear quotients are implied from $I^{2}$ to $I^{3}$, and from $I^{3}$ to $I^{4}$.
Proposition 6.5.
Assume the same setting as in Theorem 6.4. Suppose that $I^{2}$ has linear quotients with an order ${\mathcal{M}}^{(2)}$ of its generators. Then, $I^{3}$ has linear quotients with the order ${\mathcal{M}}^{(3)}$, as constructed as in Definition 6.3, of its generators.
Proof.
We proceed in the same line of arguments as that of Theorem 6.4. We need to prove that ${\mathcal{M}}^{(3)}$ gives a linear quotient order for $I^{3}$. Again, consider $A=u_{i}e_{j}$ and $B=u_{k}e_{\ell}$, where $e_{i}=ab\succ e_{\ell}=cd$, $u_{i}=L_{1}L_{2}$, $u_{k}=M_{1}M_{2}$, and $M_{t}\preceq cd$ for $t=1,2$. Case 1 follows in exactly the same way as that in Theorem 6.4.
In Case 2, if $\gamma<2$ then $Q$ is a variable and we can choose $C=A$. If $\gamma\geq 2$, then it forces $\gamma=q=2$. That means
$A$ $=(x_{1}y_{1})\cdot(x_{2}y_{2})\cdot(ab)$
$B$ $=(y_{1}z_{1})\cdot(be)\cdot(ad).$
For $Q=x_{1}x_{2}$, we must also have that $y_{2}$ is $d$ or $e$. That means, either $x_{2}d$ or $x_{2}e$ is an edge in $G$. The same argument for this situation in Theorem 6.4 now applies to complete the proof.
Proposition 6.6.
Assume the same setting as in Theorem 6.4. Suppose that $I^{3}$ has linear quotients with the order ${\mathcal{M}}^{(3)}$ of its generators. Then, so does $I^{4}$ with the order ${\mathcal{M}}^{(4)}$ of its generators.
Proof.
Again, we follow the same line of arguments as in Theorem 6.4 and, at the same time, identify the differences or where arguments in Theorem 6.4 are not applicable.
Case 1 follows with exactly the same argument. In Case 2, as noted in Proposition 6.5, we may assume that $\gamma\geq 2$. That means, either $\gamma=2$ or $\gamma=3=q$.
If $\gamma=2$, then
$A$ $=(x_{1}y_{1})\cdot(x_{2}y_{2})\cdot L_{3}\cdot(ab)$
$B$ $=(y_{1}z_{1})\cdot(y_{2}z_{2})\cdot(be)\cdot(ad).$
Since $Q=x_{1}x_{2}$, this forces $L_{3}\subseteq\{z_{1},z_{2},d,e\}$, in which case the same argument in Theorem 6.4 (when there exists $\theta$ such that $L_{\theta}\subseteq Z$) applies.
If $\gamma=3$, then
$A$ $=(x_{1}y_{1})\cdot(x_{2}y_{2})\cdot(x_{3}y_{3})\cdot(ab)$
$B$ $=(y_{1}z_{1})\cdot(y_{2}z_{2})\cdot(be)\cdot(ad),$
Since $Q=x_{1}x_{2}x_{3}$, this implies that $y_{3}$ is either $d$ or $e$. That is, either $x_{3}d$ or $x_{3}e$ is an edge in $G$. The same argument of Theorem 6.4 in this case also applies to complete the proof.
Remark 6.7.
Our proof does not currently work for the implications $I^{4}\rightarrow I^{5}$, $I^{5}\rightarrow I^{6}$ and $I^{6}\rightarrow I^{7}$.
Remark 6.8.
The proof of Theorem 6.4 applies for relatively manageable graphs to also work for the implications $I^{4}\rightarrow I^{5},I^{5}\rightarrow I^{6}$ and $I^{6}\rightarrow I^{7}$; for instance, anticycles, i.e., the complementary graphs of cycles. Particularly, by exhibiting a specific linear quotients order for $I(G)^{2}$, when $G$ is an anticycle, it follows that $I(G)^{q}$ has linear quotients for all $q\geq 2$. This is a main result of a recent preprint [4].
Acknowledgements
The research for this paper was initiated during the authors’ stay at the American Institute of Mathematics (AIM), as part of the AIM SQuaRE program. We would like to thank AIM for the warm hospitality and providing a stimulating research environment. Faridi is supported by NSERC Discovery Grant 2023-05929. Hà is partially supported by a Simons Foundation grant. Kara is supported by NSF grant DMS-2418805.